OpenStudy (anonymous):

What is the seventh term of a sequence whose first term is 1 and whose common ratio is 3? 243 729 2,187

OpenStudy (anonymous):

fan and medal to first correct awnser

OpenStudy (raden):

use this formula : Tn = a1 * r^(n-1) known a1 = , r = 3 and n = 6

OpenStudy (raden):

oppsss. n = 7

OpenStudy (anonymous):

this is first year algebra i dont know how to do that

OpenStudy (anonymous):

thanks for the explain

OpenStudy (raden):

Tn = a1 * r^(n-1) T7 = 1 * 3^(7-1) T7 = 3^6 = 729

OpenStudy (anonymous):

were u given an equation to work with dan??

OpenStudy (owlcoffee):

but, Rad, is this a Geometric sucession?

OpenStudy (anonymous):

yes i got it now @JaziBoo

OpenStudy (anonymous):

Kk, and if u need anything juz let me know ;D

OpenStudy (raden):

yeah, it is a Geometric sucession. you just have 8 terms, @Owlcoffee . the question is asking us the 7th term

OpenStudy (raden):

by manual, this sequence is 1, 3, 9, 27, 81, 243, 729 if you multiply by 3 for next term, it would be 8th term. @Owlcoffee

OpenStudy (owlcoffee):

yeah, I miscalculated, whenever you do it manually and given the first term, you always go on 1 less to the exponent (proof of the formula) Wich is easy to prove, but I'm soooo lazy...